FiND the HCF of 56 and 72# bbbnnbbnnnnnnenndnfngnnfnfnfn&nnnnmnbbnbbb

Grade:6

2 Answers

Arun
4 years ago
Dear student
 

56 = 2 2 2 7

72 = 2 2 2 3 3

HCF = 2^3 =8

 

Hope it helps

In case of any difficulty, please feel free to ask again.

 

Thanks and regards

Arun

Vikas TU
4 years ago
Dear student
Please refer the solution 
72= 56*1+16
56=16*3+8
16=8*2+0..............
Hcf is 8,,,,,,,,,,,,,,,,,,,,,.
Good Luck